what is 3-1/2 divided by 2-4/5 in simplest form
Answer by edjones(8007)

(3 - 1/2) / (2 - 4/5)
= 2 1/2 / 1 1/5
= 5/2 / 6/5
= 5/2 * 5/6
= 25/12
.
Ed